# Dynamic Contrastive Knowledge Distillation for Efficient Image Restoration

### v1

@@ Abstract
Knowledge distillation (KD) is a valuable yet challenging approach that enhances a compact student network by learning from a high-performance but cumbersome teacher model. However, previous KD methods for image restoration overlook the state of the student during the distillation, adopting a fixed solution space that limits the capability of KD. Additionally, relying solely on L1-type loss struggles to leverage the distribution information of images. In this work, we propose a novel dynamic contrastive knowledge distillation (DCKD) framework for image restoration. Specifically, we introduce dynamic contrastive regularization to perceive the student’s learning state and dynamically adjust the distilled solution space using contrastive learning. Additionally, we also propose a distribution mapping module to extract and align the pixel-level category distribution of the teacher and student models. Note that the proposed DCKD is a structure-agnostic distillation framework, which can adapt to different backbones and can be combined with methods that optimize upper-bound constraints to further enhance model performance. Extensive experiments demonstrate that DCKD significantly outperforms the state-of-the-art KD methods across various image restoration tasks and backbones.
@@ Introduction
Image restoration aims to recover high-quality images from low-quality ones degraded by processes such as sub-sampling, blurring, and rain streaks. It is a highly challenging ill-posed inverse problem since crucial content information is missing during degradation. Recently, convolutional neural networks (CNNs) (Dong et al. 2015; Lim et al. 2017; Zhang et al. 2018) and Transformers (Liang et al. 2021; Chen et al. 2021; Wang et al. 2022; Zamir et al. 2022) have been extensively investigated for designing various models, achieving remarkable success in image restoration. However, these models demand high resources and exhibit inefficiencies, making deployment on resource-constrained devices challenging (Zhang et al. 2021). To facilitate their real-world applications, there is a growing research focus on compressing image restoration models.
@@
Knowledge distillation (KD) is an effective model compression method that transfers knowledge from a cumbersome teacher model to a lightweight student model. This process allows the student models to inherit the capabilities of the teacher models, resulting in significant performance improvements while reducing computational and storage requirements. KD has gained broad recognition for its excellent performance and broad applicability. It also can be combined with other model compression techniques, such as quantization (Du et al. 2021; Ayazoglu 2021; Hong et al. 2022), pruning (Fan et al. 2020; Wang et al. 2021a; Oh et al. 2022), compact architecture design (Ahn, Kang, and Sohn 2018; Zhang et al. 2022; Chen et al. 2022a), and neural architecture search (NAS) (Gou et al. 2020; Kim et al. 2021; He et al. 2022), to enhance the compactness of student models further.
Since KD has been well-established in natural language processing (Hahn and Choi 2019; Sanh et al. 2019; Fu et al. 2021) and high-level vision tasks (Touvron et al. 2021; Lin et al. 2022; Chen et al. 2022b), researchers have been investigating KD for image restoration methods (He et al. 2020; Lee et al. 2020; Wang et al. 2021c; Zhang et al. 2023; Li et al. 2024; Jiang et al. 2024; Zhang et al. 2024). However, these methods adopt a fixed solution space, which limits their adaptability to the evolving state of the student model during the distillation process. As illustrated in Fig. 1 (a), the vanilla distillation method (Hinton, Vinyals, and Dean 2015) only constrains the upper bound of the solution space. Although existing works (Zhang et al. 2023; Jiang et al. 2024) explore more effective and diverse upper bounds, the lack of constraints on the lower bound of the output image increases the difficulty of optimizing the solution space. This often generates low-quality images with artifacts, color distortion, and blurring. As illustrated in Fig. 1 (b), CSD (Wang et al. 2021c) introduces contrastive learning to design lower-bound constraints, significantly enhancing the transfer of knowledge from the teacher. However, in the later stages of training, the student anchor moves far from the lower bound, leading to a diminished constraint effect.
@@
Overall, our main contributions are as follows:
@@
Knowledge distillation aims to significantly reduce deployment costs while improving the performance of student models by emulating the behavior of teacher models (Hinton, Vinyals, and Dean 2015; Gou et al. 2021; Lee et al. 2020). In recent years, numerous works have focused on knowledge distillation for image restoration. He et al. proposed FAKD to align the spatial affinity matrix of the feature maps between the teacher and student models (He et al. 2020).To alleviate the semantic differences between features of the teacher and student, Li et al. proposed MiPKD, which achieves feature and stochastic network block mixture in latent space (Li et al. 2024). Jiang et al. proposed MTKD that designs a composite output from multiple teachers to provide the student with a more robust teacher model (Jiang et al. 2024). However, these KD methods primarily focus on improving the upper bound of the model’s solution space, without leveraging the distribution information of the images.
@@
Recently, many researchers have explored the combination of contrastive learning and knowledge distillation to construct a comprehensive solution space (Tian, Krishnan, and Isola 2019; Xu et al. 2020; Yang et al. 2023). Wang et al. introduced contrastive learning in knowledge distillation, utilizing other images within the same dataset to construct a lower bound for the solution space (Wang et al. 2021b). Similarly, CSD (Wang et al. 2021c) proposed a contrastive self-distillation method, utilizing different images within the batch to provide lower bound constraints. Luo et al. enriched the lower bound constraints in the image deraining task by altering the direction of falling raindrops for the student model (Luo et al. 2023). However, these methods employ a fixed solution space, which leads to a weakening of the lower-bound constraint when the student anchor moves away from the lower bound in the later stages of training. Different from these methods, DCKD proposes a dynamic lower-bound constraint that progressively narrows the solution space, and can also be combined with enhanced upper-bound approaches. Additionally, DMM is proposed to extract and align the pixel-level category distribution information between the teacher and student networks.
@@
Previous KD methods (Hinton, Vinyals, and Dean 2015; Wang et al. 2021c) employed a fixed solution space, which causes the lower-bound constraints on the student model to weaken in the later stages of training. To address this problem, we propose dynamic contrastive regularization (DCR), which dynamically adjusts the solution space based on the student model’s state.
As shown in Fig. 2 (a), we first feed the input image $I_{LQ}$ into the dynamic negative sample generator, which consists of the degrading module and the history model. The degrading module applies random degrading operations on $I_{LQ}$, generating $N$ different degraded images $I_{dirty}$. The history model $\mathcal{F}_{S}^{his}$ then reconstructs these degraded images based on the historical state of the student model, producing $N$ different negative images $I_{Neg}$ as the lower-bound of the solution space:
@@
where $f^{Anc}$, $f^{Pos}$, and $f^{Neg}$ represent the features extracted from the output of the student model, positive image and negative images, respectively. $i$ denotes the $i$-th layer of Encoder. $\lambda_{i}$ is the balancing weight for $i$-th layer.
@@
Finally, we employ cross-entropy loss to align $C^{T}$ and $C^{S}$:
@@
The proposed DCKD is evaluated on three image restoration tasks, image super-resolution, image deblurring, and image draining. Following (Li et al. 2024), we verify the effectiveness of DCKD on Transformer-based SwinIR (Liang et al. 2021) and CNN-based RCAN (Zhang et al. 2018) in image super-resolution. For image deblurring, we use NAFNet (Chen et al. 2022a) and Restormer (Zamir et al. 2022) as the teacher backbones. For image deraining, we employ Restormer (Zamir et al. 2022) as the backbone. The configuration details of teacher and student models are presented in the Appendix.
Scale Method Set5 Set14 BSD100 Urban100
PSNR PSNR PSNR PSNR
$\times$2 Teacher 38.36 34.14 32.45 33.40/
Scratch 38.00 33.56 32.19 32.05
Logits 38.04 33.61 32.22 32.09
FAKD 38.03 33.63 32.21 32.06
MiPKD 38.14 33.76 32.29 32.46
DCKD 38.17 33.85 32.31 32.59
$\times$3 Teacher 34.89 30.77 29.37 29.29
Scratch 34.41 30.43 29.12 28.20
Logits 34.44 30.45 29.14 28.23
FAKD 34.42 30.42 29.12 28.18
MiPKD 34.53 30.52 29.19 28.47
DCKD 34.59 30.57 29.22 28.63
$\times$4 Teacher 32.72 28.94 27.83 27.07
Scratch 32.31 28.67 27.61 26.15
Logits 32.27 28.67 27.62 26.15
FAKD 32.22 28.65 27.61 26.09
MiPKD 32.39 28.76 27.68 26.37
DCKD 32.49 28.82 27.72 26.53
$\mathrm{\textbf{DCKD}}^{*}$ 32.51 28.88 27.74 26.62
Table 1: Quantitative comparison of the benchmark datasets for the SwinIR model. The best and second-best performances are highlighted in bold and underlined, respectively. DCKD* adds the optimized upper bound constraint.
@@
For image super-resolution, DCKD is trained using 800 images from the DIV2K dataset (Timofte et al. 2017) and evaluated on four benchmark datasets: Set5 (Bevilacqua et al. 2012), Set14 (Zeyde, Elad, and Protter 2012), BSD100 (Martin et al. 2001), and Urban100 (Huang, Singh, and Ahuja 2015). For image deblurring, the models are trained and tested both on the GoPro dataset (Nah, Hyun Kim, and Mu Lee 2017b). For image deraining, we train DCKD on 13,712 clean-rainy image pairs collected from multiple datasets (Fu et al. 2017; Yang et al. 2017; Zhang, Sindagi, and Patel 2019; Li et al. 2016) and evaluate it on Test100 (Zhang, Sindagi, and Patel 2019), Rain100H (Yang et al. 2017), Rain100L (Yang et al. 2017), Test2800 (Fu et al. 2017), and Test1200 (Zhang and Patel 2018). We employ the PSNR and SSIM (Wang et al. 2004) metrics to evaluate the restoration performance. For the image super-resolution and image deraining tasks, the metrics are computed on the Y channel in the YCbCr color space. For the image deblurring, PSNR and SSIM are evaluated in the RGB color space.
@@
For image super-resolution, the input is randomly cropped into $48\times 48$ patches and augmented by random horizontal and vertical flips and rotations. All the models are trained using ADAM optimizer (Kingma and Ba 2014) with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and $\epsilon=10^{-8}$. The training batch size is set to 16 with a total of $2.5\times 10^{5}$ iterations. The initial learning rate is set to $10^{-4}$ and is decayed by a factor of 10 at every $10^{5}$ update. DCKD is implemented by PyTorch using 4 NVIDIA V100 GPUs. For other image restoration tasks, we strictly adhere to the original training configurations of each teacher model. The training configurations for other tasks are detailed in the Appendix.
Results and Comparison
Scale Method Set5 Set14 BSD100 Urban100
PSNR PSNR PSNR PSNR
$\times$2 Teacher 38.27 34.13 32.41 33.34
Scratch 38.13 33.78 32.26 32.63
Logits 38.17 33.83 32.29 32.67
FAKD 38.17 33.83 32.29 32.65
CrossKD 38.18 33.81 32.30 32.66
MiPKD 38.21 33.92 32.32 32.83
DCKD 38.25 34.01 32.35 32.92
$\times$3 Teacher 34.74 30.65 29.32 29.09
Scratch 34.61 30.45 29.18 28.59
Logits 34.61 30.47 29.21 28.62
FAKD 34.63 30.51 29.21 28.62
CrossKD 34.65 30.50 29.21 28.60
MiPKD 34.72 30.55 29.25 28.76
DCKD 34.74 30.60 29.29 28.87
$\times$4 Teacher 32.63 28.87 27.77 26.82
Scratch 32.38 28.69 27.63 26.36
Logits 32.45 28.76 27.67 26.49
FAKD 32.46 28.77 27.68 26.50
CrossKD 32.46 28.79 27.69 26.52
MiPKD 32.46 28.77 27.69 26.55
DCKD 32.56 28.82 27.73 26.69
$\mathrm{\textbf{DCKD}}^{*}$ 32.58 28.86 27.74 26.74
Table 2: Quantitative comparison of the benchmark datasets for the RCAN model. The best and second-best performances are highlighted in bold and underlined, respectively. DCKD* adds the optimized upper bound constraint.
@@
Table 3: Quantitative comparison on the GoPro dataset for the deblurring task.
Figure 4: The image deblurring visual comparison of NAFNet on the GoPro dataset.
@@
We compare our framework with the representative KD methods: train from scratch, Logits (Hinton, Vinyals, and Dean 2015), FAKD (He et al. 2020), CrossKD (Fang et al. 2023), and MiPKD (Li et al. 2024), on ×$\times 2$, ×$\times 3$, and ×$\times 4$ super-resolving scales.
The quantitative results for transformer-based backbone SwinIR and CNN-based backbone RCAN are presented in Tab. 1 and Tab. 2, respectively. Existing KD methods provide limited improvement for the student model, and in some cases, they even lead to worse performance compared to models trained without KD. For example, using FAKD for distillation on Urban100 results in poorer performance than training the model from scratch on SwinIR. DCKD is effective for both Transformer-based and CNN-based architectures, significantly outperforming existing KD methods by more than 0.1dB across all three scales on the Urban100 dataset. We deliberately used the most straightforward approach to demonstrate DCKD, showing that dynamic lower-bound constraints can yield strong results even without improving upper-bound constraints. To demonstrate that DCKD can be combined with methods that optimize upper-bound constraints, we incorporated DUKD (Zhang et al. 2023) into the DCKD framework to enhance upper-bound constraints, resulting in DCKD*. As we can see, the proposed DCKD can be combined with existing KD methods that optimize the upper bound, further significantly enhancing performance. DCKD* significantly outperforms the SOTA method MiPKD by 0.25dB on SwinIR and 0.19dB on RCAN at ×$\times 4$.
Figure 5: The ×$\times 4$ super-resolution visual comparison of SwinIR on the Urban100 dataset.
Figure 6: The ×$\times 4$ super-resolution visual comparison of RCAN on the Urban100 dataset.
@@
Scratch 31.01/0.9122 30.51/0.8932 37.47/0.9714 33.78/0.9396 33.67/0.9295 -
Logits 31.04/0.9143 30.48/0.8915 37.17/0.9712 33.81/0.9399 33.78/0.9310 -
Our 31.08/0.9167 30.54/0.8969 38.02/0.9762 33.91/0.9411 33.95/0.9326 3.8M
Table 4: Quantitative comparison of the benchmark datasets for the deraining task.
Fig. 5 and Fig. 6 present challenging visual examples for Transformer and CNN backbones, respectively. Compared to existing KD methods, our approach enables the student model to better capture structural textures, such as more accurately reconstructing sidewalk lines and building structures. More visual comparisons for various examples and models are presented in the Appendix.
@@
Tab. 3 provides the quantitative comparison on the GoPro dataset (Nah, Hyun Kim, and Mu Lee 2017b). Our method demonstrates consistent effectiveness on both the CNN-based NAFNet (Chen et al. 2022a) and the Transformer-based Restormer (Zamir et al. 2022). Compared to the Logits KD, our DCKD achieves 0.17dB performance improvement on different backbones. Moreover, with comparable parameters, the student model of NAFNet significantly outperforms MT-RNN (Park et al. 2020) by 0.28dB. Fig. 4 illustrates the deblurring visualization results. DCKD restores the clearest window outlines, significantly enhancing the deblurring capability of the student model.
Figure 7: The image deraining visual comparison of Restormer on the Rain100L dataset.
@@
Tab. 4 shows the performance of different methods on several benchmark datasets for image draining. With only 18.9% of MPRNet’s parameters, DCKD significantly surpasses it by 1.8dB in PSNR on the Rain100L dataset. Additionally, compared to other distillation methods, DCKD outperforms Logits KD by 0.55dB on the Rain100L. Fig. 7 presents a visual comparison of derained images. DCKD further enhances the student’s ability to remove rain streaks compared to the logit distillation method.
@@
Table 6: Ablation study on the type of degradation.
@@
Table 7: Ablation study on the balancing weights of our framework.
@@
For ablation experiments, we train DCKD on RCAN (Zhang et al. 2018) with the ×$\times 2$ scale for the SR task. We then validate the results on the Set14 (Zeyde, Elad, and Protter 2012) and Urban100 (Huang, Singh, and Ahuja 2015).
@@
As Tab. 5, we first conduct an ablation study on the two main modules of DCKD. The results show that our DCR and DMM outperform the baseline by 0.16dB and 0.14dB on the Urban100, respectively. This demonstrates the effectiveness of the proposed modules. Additionally, the combination of DCR and DMM further enhances model performance, achieving improvements of 0.18dB on the Set14 and 0.25dB on the Urban100 compared to the baseline.
Type of Random Degradation
We conduct ablation studies on the type of degradation module in the Dynamic Negative Sample Generator (DNSG), including random blurring, random resize, random noise, and mixed degradation. As shown in Tab. 6, random noise achieves the best results, while random blurring and random resize produce a similar performance. The mixed degradation results in lower-quality negative samples, which weakens the lower bound constraint and leads to the worst performance. Consequently, we adopted random noise degradation in DCKD.
Impact of Balancing Weights
@@
The impact of the number of negative samples is reported in Fig. 8 (a). The results indicate that as the number of negative samples increases, the performance improves consistently. However, when the number of negative samples exceeds 5, the performance gains diminish while significantly increasing training time and memory costs. Therefore, we set the number of negative samples to 5.
Impact of Initial Update Step

### v2

@@ Abstract
Knowledge distillation (KD) is a valuable yet challenging approach that enhances a compact student network by learning from a high-performance but cumbersome teacher model. However, previous KD methods for image restoration overlook the state of the student during the distillation, adopting a fixed solution space that limits the capability of KD. Additionally, relying solely on L1-type loss struggles to leverage the distribution information of images. In this work, we propose a novel dynamic contrastive knowledge distillation (DCKD) framework for image restoration. Specifically, we introduce dynamic contrastive regularization to perceive the student’s learning state and dynamically adjust the distilled solution space using contrastive learning. Additionally, we also propose a distribution mapping module to extract and align the pixel-level category distribution of the teacher and student models. Note that the proposed DCKD is a structure-agnostic distillation framework, which can adapt to different backbones and can be combined with methods that optimize upper-bound constraints to further enhance model performance. Extensive experiments demonstrate that DCKD significantly outperforms the state-of-the-art KD methods across various image restoration tasks and backbones. Our codes are available at https://github.com/super-SSS/DCKD.
@@ Introduction
Image restoration aims to recover high-quality images from low-quality ones degraded by processes such as sub-sampling, blurring, and rain streaks. It is a highly challenging ill-posed inverse problem since crucial content information is missing during degradation. Recently, convolutional neural networks (CNNs) (Dong et al. 2015; Lim et al. 2017; Zhang et al. 2018) and Transformers (Liang et al. 2021; Chen et al. 2021; Wang et al. 2022; Zamir et al. 2022) have been extensively investigated for designing various models, achieving remarkable success in image restoration. However, these models demand high resources and exhibit inefficiencies, making deployment on resource-constrained devices challenging. To facilitate their real-world applications, there is a growing research focus on compressing image restoration models.
@@
Knowledge distillation (KD) is an effective model compression method that transfers knowledge from a cumbersome teacher model to a lightweight student model. This process allows the student model to inherit the capabilities of the teacher model, resulting in significant performance improvements while reducing computational and storage requirements. KD has gained broad recognition for its excellent performance and broad applicability. It also can be combined with other model compression techniques, such as quantization (Du et al. 2021; Ayazoglu 2021; Hong et al. 2022), pruning (Fan et al. 2020; Wang et al. 2021a; Oh et al. 2022), compact architecture design (Ahn, Kang, and Sohn 2018; Zhang et al. 2022; Chen et al. 2022a), and neural architecture search (NAS) (Gou et al. 2020; Kim et al. 2021; He et al. 2022), to enhance the compactness of student models further.
Since KD has been well-established in natural language processing (Hahn and Choi 2019; Sanh et al. 2019; Fu et al. 2021) and high-level vision tasks (Touvron et al. 2021; Lin et al. 2022; Chen et al. 2022b), researchers have been investigating KD for image restoration methods (He et al. 2020; Lee et al. 2020; Wang et al. 2021d; Zhang et al. 2023; Li et al. 2024; Jiang et al. 2024; Zhang et al. 2024). However, these methods adopt a fixed solution space, which limits their adaptability to the evolving state of the student model during the distillation process. As illustrated in Fig. 1 (a), the vanilla distillation method (Hinton, Vinyals, and Dean 2015) only constrains the upper bound of the solution space. Although existing works (Zhang et al. 2023; Jiang et al. 2024) explore more effective and diverse upper bounds, the lack of constraints on the lower bound of the output image increases the difficulty of optimizing the solution space. This often generates low-quality images with artifacts, color distortion, and blurring. As illustrated in Fig. 1 (b), CSD (Wang et al. 2021d) introduces contrastive learning to design lower-bound constraints, significantly enhancing the transfer of knowledge from the teacher. However, in the later stages of training, the student anchor moves far from the lower bound, leading to a diminished constraint effect.
@@
Overall, our main contributions are summarized as:
@@
Knowledge distillation aims to significantly reduce deployment costs while improving the performance of student models by emulating the behavior of teacher models (Hinton, Vinyals, and Dean 2015; Lee et al. 2020; Gou et al. 2021). In recent years, numerous works have focused on knowledge distillation for image restoration. He et al. proposed FAKD to align the spatial affinity matrix of the feature maps between the teacher and student models (He et al. 2020). To alleviate the semantic differences between features of the teacher and student, Li et al. proposed MiPKD, which achieves feature and stochastic network block mixture in latent space (Li et al. 2024). Jiang et al. proposed MTKD that designs a composite output from multiple teachers to provide the student with a more robust teacher model (Jiang et al. 2024). However, these KD methods primarily focus on improving the upper bound of the model’s solution space, without leveraging the distribution information of the images.
@@
Recently, many researchers have explored the combination of contrastive learning and knowledge distillation to construct a comprehensive solution space (Tian, Krishnan, and Isola 2019; Xu et al. 2020; Yang et al. 2023). Wang et al. introduced contrastive learning in knowledge distillation, utilizing other images within the same dataset to construct a lower bound for the solution space (Wang et al. 2021b). Similarly, CSD (Wang et al. 2021d) proposed a contrastive self-distillation method, utilizing different images within the batch to provide lower bound constraints. Luo et al. enriched the lower bound constraints in the image deraining task by altering the direction of falling raindrops for the student model (Luo et al. 2023). However, these methods employ a fixed solution space, which leads to a weakening of the lower-bound constraint when the student anchor moves away from the lower bound in the later stages of training. Different from these methods, DCKD proposes a dynamic lower-bound constraint that progressively narrows the solution space, and can also be combined with enhanced upper-bound approaches. Additionally, DMM is proposed to extract and align the pixel-level category distribution information between the teacher and student networks.
@@
Previous KD methods (Hinton, Vinyals, and Dean 2015; Wang et al. 2021d) employed a fixed solution space, which causes the lower-bound constraints on the student model to weaken in the later stages of training. To address this problem, we propose dynamic contrastive regularization (DCR), which dynamically adjusts the solution space based on the student model’s state.
As shown in Fig. 2 (a), we first feed the input image $I_{LQ}$ into the dynamic negative sample generator, which consists of the degradation module and the history model. The degradation module applies random degradation operations on $I_{LQ}$, generating $N$ different degraded images $I_{dirty}$. The history model $\mathcal{F}_{S}^{his}$ then reconstructs these degraded images based on the historical state of the student model, producing $N$ different negative images $I_{Neg}$ as the lower-bound of the solution space:
@@
where $f^{Anc}$, $f^{Pos}$, and $f^{Neg}$ represent the features extracted from the output of the student model, positive image and negative images, respectively. $i$ denotes the $i$-th layer of Encoder. $\lambda_{i}$ is the balancing weight for the $i$-th layer.
@@
Finally, we use cross-entropy loss to align $C^{T}$ and $C^{S}$:
@@
The proposed DCKD is evaluated on three image restoration tasks: image super-resolution, image deblurring, and image deraining. Following (Li et al. 2024), we verify the effectiveness of DCKD on Transformer-based SwinIR (Liang et al. 2021) and CNN-based RCAN (Zhang et al. 2018) in image super-resolution. For image deblurring, we use NAFNet (Chen et al. 2022a) and Restormer (Zamir et al. 2022) as the teacher backbones. For image deraining, we employ Restormer as the backbone. The configuration details of teacher and student models are presented in Tab. 1.
Model Role Channel Block/Group #Params
SwinIR Teacher 180 6/- 11.9M
Student 60 4/- 1.2M
RCAN Teacher 64 20/10 15.6M
Student 64 6/10 5.2M
NAFNet Teacher 32 36/- 17.1M
Student 16 22/- 2.7M
Restormer Teacher 48 44/- 26.1M
Student 24 22/- 3.8M
Table 1: The specifications of teacher and student models.
@@
For image super-resolution, DCKD is trained using 800 images from DIV2K (Timofte et al. 2017) and evaluated on four benchmark datasets. For image deblurring, the models are trained and tested both on GoPro dataset (Nah, Hyun Kim, and Mu Lee 2017b). For image deraining, we train DCKD on 13,712 clean-rainy image pairs collected from multiple datasets (Fu et al. 2017; Yang et al. 2017; Zhang, Sindagi, and Patel 2019; Li et al. 2016) and evaluate it on Test100 (Zhang, Sindagi, and Patel 2019), Rain100H (Yang et al. 2017), Rain100L (Yang et al. 2017), Test2800 (Fu et al. 2017), and Test1200 (Zhang and Patel 2018). We employ the PSNR and SSIM (Wang et al. 2004) metrics to evaluate the restoration performance. For image super-resolution and image deraining tasks, the metrics are computed on the Y channel in the YCbCr color space. For image deblurring, PSNR and SSIM are evaluated in the RGB color space.
@@
For image super-resolution, the input is randomly cropped into $48\times 48$ patches and augmented by random horizontal and vertical flips and rotations. All the models are trained using ADAM optimizer (Kingma and Ba 2014) with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and $\epsilon=10^{-8}$. The training batch size is set to 16 with a total of $2.5\times 10^{5}$ iterations. The initial learning rate is set to $10^{-4}$ and is decayed by a factor of 10 at every $10^{5}$ update. DCKD is implemented by PyTorch using 4 NVIDIA V100 GPUs. For other image restoration tasks, we strictly adhere to the original training configurations of each teacher model. More training configurations for other tasks are presented in the Appendix [^2] https://arxiv.org/abs/2412.08939.
Scale Method Model Set5 Set14 BSD100 Urban100 Model Set5 Set14 BSD100 Urban100
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
$\times$2 Teacher SwinIR 38.36/0.9620 34.14/0.9227 32.45/0.9030 33.40/0.9394 RCAN 38.27/0.9614 34.13/0.9216 32.41/0.9027 33.34/0.9384
Scratch 38.00/0.9607 33.56/0.9178 32.19/0.9000 32.05/0.9279 38.13/0.9610 33.78/0.9194 32.26/0.9007 32.63/0.9327
Logits 38.04/0.9608 33.61/0.9184 32.22/0.9003 32.09/0.9282 38.17/0.9611 33.83/0.9197 32.29/0.9010 32.67/0.9329
FAKD 38.03/0.9608 33.63/0.9182 32.21/0.9001 32.06/0.9279 38.17/0.9612 33.83/0.9199 32.29/0.9011 32.65/0.9330
MiPKD 38.14/0.9611 33.76/0.9194 32.29/0.9011 32.46/0.9313 38.21/0.9613 33.92/0.9203 32.32/0.9015 32.83/0.9344
DCKD 38.17/0.9613 33.85/0.9205 32.31/0.9017 32.59/0.9330 38.25/0.9615 34.01/0.9213 32.35/0.9019 32.92/0.9351
$\times$3 Teacher 34.89/0.9312 30.77/0.8503 29.37/0.8124 29.29/0.8744 34.74/0.9299 30.65/0.8482 29.32/0.8111 29.09/0.8702
Scratch 34.41/0.9273 30.43/0.8437 29.12/0.8062 28.20/0.8537 34.61/0.9288 30.45/0.8444 29.18/0.8074 28.59/0.8610
Logits 34.44/0.9275 30.45/0.8443 29.14/0.8066 28.23/0.8545 34.61/0.9291 30.47/0.8447 29.21/0.8080 28.62/0.8612
FAKD 34.42/0.9273 30.42/0.8437 29.12/0.8062 28.18/0.8533 34.63/0.9290 30.51/0.8453 29.21/0.8079 28.62/0.8612
MiPKD 34.53/0.9283 30.52/0.8456 29.19/0.8079 28.47/0.8591 34.72/0.9296 30.55/0.8458 29.25/0.8087 28.76/0.8640
DCKD 34.59/0.9291 30.57/0.8475 29.22/0.8093 28.63/0.8633 34.74/0.9299 30.60/0.8472 29.29/0.8099 28.87/0.8662
$\times$4 Teacher 32.72/0.9021 28.94/0.7914 27.83/0.7459 27.07/0.8164 32.63/0.9002 28.87/0.7889 27.77/0.7436 26.82/0.8087
Scratch 32.31/0.8955 28.67/0.7833 27.61/0.7379 26.15/0.7884 32.38/0.8971 28.69/0.7842 27.63/0.7379 26.36/0.7947
Logits 32.27/0.8954 28.67/0.7833 27.62/0.7380 26.15/0.7887 32.45/0.8980 28.76/0.7860 27.67/0.7400 26.49/0.7982
FAKD 32.22/0.8950 28.65/0.7831 27.61/0.7380 26.09/0.7870 32.46/0.8980 28.77/0.7860 27.68/0.7400 26.50/0.7980
MiPKD 32.39/0.8971 28.76/0.7854 27.68/0.7403 26.37/0.7956 32.46/0.8982 28.77/0.7860 27.69/0.7402 26.55/0.7998
DCKD 32.49/0.8991 28.82/0.7877 27.72/0.7422 26.53/0.8007 32.56/0.8995 28.82/0.7877 27.73/0.7423 26.69/0.8041
$\mathrm{\textbf{DCKD}}^{*}$ 32.51/0.8992 28.88/0.7890 27.74/0.7430 26.62/0.8032 32.58/0.8996 28.86/0.7885 27.74/0.7425 26.74/0.8054
Table 2: Quantitative comparison on the benchmark datasets for image super-resolution. The best and second-best performances are highlighted in bold and underlined, respectively. The FAKD results on SwinIR are from our reproduction experiments.
@@
Table 3: Quantitative comparison for image deblurring.
Figure 4: Visual comparison for image deblurring.
Results and Comparison
@@
We compare our framework with the representative KD methods: train from scratch, Logits (Hinton, Vinyals, and Dean 2015), FAKD (He et al. 2020), and MiPKD (Li et al. 2024), on ×$\times 2$, ×$\times 3$, and ×$\times 4$ super-resolving scales.
The quantitative results for SwinIR and RCAN are presented in Tab. 2. Existing KD methods provide limited improvement for the student models, and in some cases, they even lead to worse performance compared to models trained without KD. For example, using FAKD for distillation on Urban100 results in worse performance than training the model from scratch on SwinIR. DCKD is effective for both Transformer-based and CNN-based architectures, significantly outperforming existing KD methods by more than 0.1dB across all three scales on Urban100. We deliberately use the most straightforward approach to demonstrate DCKD, showing that dynamic lower-bound constraints can yield strong results even without improving upper-bound constraints. To demonstrate that DCKD can be combined with methods that optimize upper-bound constraints, we incorporate DUKD (Zhang et al. 2023) into the DCKD framework to enhance upper-bound constraints, resulting in DCKD*. As we can see, the proposed DCKD can be combined with existing KD methods that optimize the upper bound, further significantly enhancing performance. DCKD* significantly outperforms the SOTA method MiPKD by 0.25dB on SwinIR and 0.19dB on RCAN at ×$\times 4$.
Fig. 5 and Fig. 6 present challenging visual examples for Transformer and CNN backbones, respectively. Compared to existing KD methods, our approach enables the student models to better capture structural textures, such as more accurately reconstructing sidewalk lines and building structures. More visual comparisons for various examples and models are presented in the Appendix.
Figure 5: The visual comparison of distilling SwinIR on Urban100 for ×$\times 4$ SR.
Figure 6: The visual comparison of distilling RCAN on Urban100 for ×$\times 4$ SR.
@@
Scratch 31.01/0.9122 30.51/0.8932 37.47/0.9714 33.78/0.9396 33.67/0.9295 3.8M
Logits 31.04/0.9143 30.48/0.8915 37.17/0.9712 33.81/0.9399 33.78/0.9310 3.8M
DCKD 31.08/0.9167 30.54/0.8969 38.02/0.9762 33.91/0.9411 33.95/0.9326 3.8M
Table 4: Quantitative comparison on the benchmark datasets for image deraining.
@@
Tab. 3 provides a quantitative comparison on GoPro dataset. Our method demonstrates consistent effectiveness across both CNN-based NAFNet and Transformer-based Restormer. Compared to the Logits KD, our DCKD achieves 0.17dB improvement on different backbones. Moreover, with comparable parameters, the student model of NAFNet significantly outperforms MT-RNN (Park et al. 2020) by 0.28dB. Fig. 4 illustrates the deblurring visualization results. DCKD restores the clearest window outlines, significantly enhancing the deblurring capability of the student model.
Figure 7: Visual comparison for image deraining.
@@
Tab. 4 shows the performance of various methods on several benchmark datasets for image deraining. With only 18.9% of MPRNet’s (Zamir et al. 2021) parameters, DCKD significantly surpasses it by 1.8dB on Rain100L dataset. Additionally, compared to other distillation methods, DCKD outperforms the Logits KD by 0.55dB on Rain100L. Fig. 7 provides a visual comparison of derained images. DCKD further enhances the student’s ability to remove rain streaks compared to the logits distillation method.
@@
Table 6: Ablation study on the degradation module.
@@
Table 7: Ablation study on the balancing weights.
@@
For ablation experiments, we train DCKD on RCAN for the SR task with a scaling factor of ×$\times 2$. We then validate the results on Set14 and Urban100 datasets.
@@
As shown in Tab. 5, we first conduct an ablation study on the two main modules of DCKD. The results indicate that our DCR and DMM outperform the baseline by 0.16dB and 0.14dB in PSNR on Urban100, respectively. This demonstrates the effectiveness of the proposed modules. Furthermore, the combination of DCR and DMM further enhances model performance, achieving PSNR improvements of 0.18dB on Set14 and 0.25dB on Urban100 compared to the baseline.
Impact of the Degradation Module
We conduct an ablation study on the impact of the degradation module in the Dynamic Negative Sample Generator (DNSG). Following Real-ESRGAN (Wang et al. 2021c), the degradation module is simple to implement by adding the Gaussian blur, Gaussian noise, resize operation (i.e., downsampling and then upsampling), or mixed degradation. The degradation results are summarized in Tab. 5 and Tab. 6. We observe that the addition degradation module (in Tab. 6) achieves at least 0.06dB PSNR improvement on Urban100, compared to that without degradation (only DMM in Tab. 5). Furthermore, the degradation with random noise achieves the highest PSNR of 32.92dB, which outperforms randomly mixed degradation by 0.05dB PSNR.
Impact of the Balancing Weights
@@
The impact of the number of negative samples is reported in Fig. 8 (a). The results indicate that as the number of negative samples increases, the performance improves consistently. However, when the number of negative samples exceeds 5, the performance gains diminish while significantly increasing training time and memory costs. Therefore, the number of negative samples is set to 5, which achieves the best trade-off between PSNR and training time.
Impact of the Initial Update Step